Imaging
TOP
(iTOP) Detector Readout

Matt Andrew,
Matt Barrett, Tom Browder, Shawn Dubey, Michael Jones, Brian
Kirby,
Isar Mostafanezhad, Luca
Macchiarulo, Marc Rosen, Xiaowen
Shi, Gary Varner
KEK TOP Module Operations: [link]
iTOP Calibration Master Module: [link]
Most current WBS 1.3
Costing (JAN-2015)
Scheduling:
- Board-stack testing/firmware/software details
documentation [link]
- Calibration Master Module [link]
- Possible beam-test configurations [pdf]
- SLAC T-518 ESTB beam test page
- WBS MIA items [link]
iTOP Previous Hot Items
(e.g. CD-1, CDR, Peer Review; FNAL 2011 beam test)
Fuji
Hall (equip list, schedule, mechanics)
[Hayakawa-san's modified version] [TWiki
page] [CRT
Elog]
F2
layout Diagram [pdf]
LEPS IRS3B Data Analysis Task list [link]
LEPS/local analysis: [twiki
link] [iTOP/IRS3 Elog]
IRS3B (w/LEPS data) available: [link]
HV/Front board links (from Gerard): [HVB
link] [Front
link]
Mechatronics Reference (30-SEP-2013) [link]
Working Group Meeting @ SLAC (28-SEP-2013) [link]
Weekly
meetings now
being archived
on
Indico
(rest kept
here as
repository)
15-APR-2014 --
Local
iTOP meeting
(Tuesday @
2pm)
To allow easier upload of files, agenda/items to show have been
moved for recent meetings to physics Indico page.
4-MAR-2014
--
Local
iTOP meeting
(Tuesday @
2pm)
- CD-2/3
Presentation Preparations (Gary/Brian/Tom)
- Posted Introductory talk? (Tom)
- Revised slides needed for Session 1A-2 (Matt B,
all testing)
- Simple KLM readout plot for Session 1B-1? (Brian)
- Intermediate board stack result for Session 1B-2
(all testing)
- Follow-ups from last week
- HV ordering status? (Brian)
- Revised laser test set-up (Brian/Matt A/Shawn)
- Assembly fixture, other mechanicals update (Marc)
- Check-lists to post (work from):
- prior to qualifying Intermediate, integrated board
stacks
- prior to 2x4 PMT scanning
- for CRT operation (need to re-machine bar-box
extension for FNAL prototype)
- Initial discussions on T-518 preparation -- ball still
in Gary's court
- Resource request survey (need to work)
- Proposal to use 1" baseplate with
- Beam height/table configuration
- Requirement: "vertical" test configuration for
cosmics
- Analysis Updates ? (all)
- AOB?
25-FEB-2014
--
Local
iTOP meeting
(Tuesday @
2pm)
- Testing Schedule
Update (Matt A)
- Side spacers from Cincinnati (how many?) [these
are production only? -- needed items for Intermed
testing/CRT?] (Matt A)
- HV ordering status? (Brian)
- Revised laser test set-up (Brian/Matt A/Shawn)
- Assembly fixture update (Marc)
- Check-lists to post (work from):
- prior to qualifying Intermediate, integrated board
stacks
- prior to 2x4 PMT scanning
- for CRT operation (need to re-machine bar-box
extension for FNAL prototype)
- Initial discussions on T-518 preparation
- Resource request survey (need to work)
- Proposal to use 1" baseplate with
- Beam height/table configuration
- Requirement: "vertical" test configuration for
cosmics
- Analysis Updates ? (all)
- What to present at tomorrow's bPID meeting?
- Updated schedule --> Matt A can you show (with
latest from Gerard) a 1 pager? High level points only:
- when board stacks ready (c13 Rev. D back yet?)
- when assy jigs ready
- when HV ready to support item 4.2 above?
- Progress on Nhit matching? (Matt B)
- AOB?
18-FEB-2014
--
Local
iTOP meeting
(Tuesday @
2pm)
- Testing Schedule
Update (Brian/Matt A)
- HV ordering status? (Brian)
- Revised laser test set-up (Brian/Matt A/Shawn)
- Assembly fixture [pdf]
discussion (Marc)
- Check-list prior to qualifying Intermediate, integrated
board stacks
- Check-list prior to 2x4 PMT scanning
- Check-list for CRT operation (need to re-machine
bar-box extension for FNAL prototype)
- Initial discussions on T-518 preparation
- Location of rotating stage
- Beam height/table configuration
- "vertical" test configuration for cosmics
- Analysis Updates ? (all)
- What to present at tomorrow's bPID meeting?
- Updated schedule?
- Status of Intermediate board stack?
- AOB?
21-JAN-2014
--
Local
iTOP meeting
(Tuesday @
2pm)
- Next stage testing Schedule
Update (Brian/Matt A)
- HV ordering status?
- Revised laser test set-up
- Check-list prior to qualifying Intermediate, integrated
board stacks
- Check-list prior to 2x4 PMT scanning
- Check-list for CRT operation (need to re-machine
bar-box extension for FNAL prototype)
- Electro-mechanical Board-stack Update (Marc/Matt)
- Action items from mini-Review/for upcoming BPAC
(Gary)
- Quantifying effectiveness of timebase feedback [data in
hand]
- "Recovering" saturated pulse data and demonstration of
timing thus achieved
- Codify a set of ASIC, Carrier, SCROD and full
board-stack test/qualification procedures [needed]
- Analysis Updates ? (all)
- What to present at tomorrow's bPID meeting?
- Updated schedule?
- Status of Intermediate board stack?
- AOB?
7-JAN-2014
--
Local
iTOP meeting
(Tuesday @
2pm)
- Hardware/firmware/test Schedule
Update (Brian/Matt A)
- Electro-mechanical Board-stack Update (Marc/Matt)
- Action items from mini-Review and beyond (Gary,
all)
- Quantifying effectiveness of timebase feedback [data in
hand]
- "Recovering" saturated pulse data and demonstration of
timing thus achieved
- Fine position (narrow laser beam) gain scan [variable
amps needed?]
- Codify a set of ASIC, Carrier, SCROD and full
board-stack test/qualification procedures
- Director's Review Presentation preparations (all)
- Status of WBS1.03 Cost Book
- Viability of IRS3C for initial Production ?
- Analysis Updates ? (all)
- AOB?
17-DEC-2013
--
Local
iTOP meeting
(Tuesday @
2pm)
- Hardware/firmware/test Schedule
Update (Brian/Matt A)
- Electro-mechanical Board-stack Update (Marc/Matt)
- Data-based feedback update (Bostjan)
- Plans for "production" board stack testing and task
sharing (Brian)
- Brief status update on software/MC (Matt B)
- Director's Review Presentation preparations (all)
- AOB?
26-NOV-2013
--
Local
iTOP meeting
(Tuesday @
2pm)
- Hardware/firmware/test Schedule
Update (Brian/Bostjan/Matt A)
- Electro-mechanical (machining/parts assy for Mini-Review)
Update (Marc/Matt)
- Mini-Review Presentation preparations (all)
- data/MC comparisons (Mike/Matt B)
- beam test overview (Gary)
- Electronics testing/status (Brian)
- AOB?
5-NOV-2013
--
Local
iTOP meeting
(Tuesday @
2pm)
- Hardware/firmware/test Schedule
Update (Brian/Bostjan/Matt A)
- Status of testing (Bostjan, Brian, all)
- "production" Rev. C testing
- additional HV
- laser test station plans
- Mechanics (machining) Update (Marc)
- Boardstack Update (Matt A carrier
revD; SCROD
fake revB) [mechatronics
reference]
- Analysis updates/plans (all)
- Discussion on what to present tomorrow (Gary, all)
- AOB?
15-OCT-2013
--
Local
iTOP meeting
(Tuesday @
2pm)
- Hardware/firmware/test Schedule
Update (Brian/Bostjan/Matt A)
- Additional HV, mechanics for next round laser
testing (Brian, all)
- Mechanics (machining) Update (Marc)
- Boardstack
Update (Matt A) [reference
link]
- Analysis updates/plans (all)
- Discussion on what to present
tomorrow (Gary, all)
- AOB?
8-OCT-2013
--
Local
iTOP meeting
(Tuesday @
2pm)
- Schedule Update (all)
- Cooling issues for laser testing (Marc, all)
- Mechanics (machining) follow-up (Marc)
- Boardstack plans (Matt A)
- Analysis updates/plans? (all)
- Discussion on what to present tomorrow (Gary, all)
- AOB?
1-OCT-2013
--
Local
iTOP meeting
(Tuesday @
2pm)
- Review Material from Reviews at SLAC
(Gary)
- IRSX
Review, Mechatronics & Working group meeting
[above]
- Mechanics (machining) follow-up (Marc)
- Short-term boardstack plans (Matt A)
- Draft
Schedule Presentation [initially for Wednesday, to be
refined for EB] (Gary]
30-SEP-2013
--
Global/Local
iTOP analysis
meeting
(Monday @ 2pm)
- Global LEPS analysis call-in (Matt B lead)
- Brief software meeting summary, status of MC for beam
test and CRT (Matt B)
- Testing/Analysis updates (round-table)
- Next testing plans (Carrier02, Rev. C)
(Bostjan/Gary)
- AOB ?
24-SEP-2013
--
Local
iTOP meeting
(Tuesday @
2pm)
- Review Information on Engineering Issues
(Gary)
- Mechanics (machining) follow-up (Marc)
- Testing/Analysis updates (round-table)
- IRSX Review preparations (Gary)
- AOB ?
20-SEP-2013
--
Engineering
Issues
follow-ups
- MCP-PMT related:
- [Summary] of
pin length, HV discharge issues (Hirose-san)
- HV/front board electro-mechanics (PNNL for Gerard)
- End-view of HV + PMT assy [png]
- Front Board contact layout (pogo-pin lands) [pdf]
- SL10-HV board [pdf]
- MOCKUP PMT unit assy [pdf]
- front-front.rev D [pdf]
- Board stack
- Rev C carrier-based [pdf]
17-SEP-2013
--
Local
iTOP meeting
(Tuesday @
2pm)
- Mechanics for Rev C
board stack machining/cooling [Gary/Marc]
- Revised fiber transceiver/cable routing diagrams [pdf]
- What to present tomorrow?
- Schedule update [for discussion Wednesday]
- Design Reviews @ SLAC [Sept 27, 28]
- AOB?
3-SEP-2013
--
Local
iTOP meeting
(Tuesday @
2pm)
- BPAC presentation preparations (all)
- Any analysis updates [new]?
- Revised board
schedule (when next CRT/beam test ops?)
- Pointers to getting started with beamtest MC in basf2,
including beta resolution
(Marko S.)
- Table to find better optical matching compounds for PMT
coupling [xls]
(Jim F.)
- AOB?
20-AUG-2013
--
Local
iTOP meeting
(Tuesday @
2pm)
- Improved Dark Box/Mechanics/cooling for next stage
testing (all)
- Any analysis updates [new]?
- Revised Analysis Schedule (Brian)
- Coordinating local analysis efforts, follow-up:
- Status/plans for efficiency scans ?
(Bostjan/Brian)
- Timing studies remaining board stacks (Brian)
- "High Current Event" testing
discussion (all)
- Coordinating shifts to muscle through
measurements (all?)
- Timing/efficiency measurements modified boards
(few channels?)
- Timing/efficiency studies for c02 Rev. C (boards
back "next week")
- What to present bPID tomorrow? (all)
- Revised Fabrication Schedule/WBS (Matt A/Gary)
- AOB ?
23-JUL-2013
--
Local
iTOP meeting
(Tuesday @
2pm)
- Review of [Action
Items] from Dedicated E-O-M meeting @ VT
3-JUL-2013
- Reminder: dedicated analysis meeting (Monday @
2pm) [updates here]
- Agenda items for Wednesday:
- Timing updates
- Efficiency updates
16-JUL-2013
--
Local
iTOP meeting
- Dedicated analysis meeting (Monday @ 2pm) Laser setup
tests
- Agenda to be updated subsequently
24-JUN-2013
/ 1-JUL-2013
--
iTOP
meeting
(deferred)
- Focus on data analysis/laser scan set-up (all)
- B2GM preparations
17-JUN-2013
--
iTOP
meeting
(2-3pm)
- Preparations for B2GM/Dedicated meeting (all)
- Electro-Mechanical-Optical Dedicated meeting (July 3rd)
- iTOP beamtest data analysis (details below)
- Cabling/GND review (outside review)
- Local preparation for laser timing data for board
stacks (Bostjan, Brian, Matt A)
- Hirose-san's visit next week (Jun 24 - July 1)
- Brief summary of Software meeting (Matt B)
- Analysis
plans/coverage discussion/coordination (Brian --
updated)
- Input to the TDR/other documents (all -- defer at
this point?)
- AOB?
29-APR-2013
--
iTOP
meeting
(2-3pm)
- Update from Fuji Hall (Brian)
- Debugging strategy (inclusive of Torino, Nagoya
collegue contributions)
- Manpower/schedule
discussion
- Decision on Integration schedule
- Brief summary of Software meeting (Matt B)
- CRT, other data analysis (Mike, all)
- Status of data taking in Hawaii, plans (Matt A,
all)
- LEPS laser analysis
update [txt] (Mike)
- LEPS (SPring-8) beam test June 1 [4] - 14 [schedule
page] [details
on fill modes]
- AOB?
15-APR-2013
--
iTOP
meeting
(2-3pm)
- Call into Gemba (Brian/Matt A)
- Brief summary of Software meeting (Matt B)
- CRT data analysis (Mike)
- Discussion on priorities (all)
- LEPS (SPring-8) beam test May 9 - 18 [schedule
page] [details
on fill modes]
- Data locations:
- panko.phys.hawaii.edu:/data3/crt_data/20130409_irs3BTests/scrod32
- cc.kek.jp:/group/belle/users/craigb/crt/backend-debug
- olympus.pnl.gov:/pic/projects/belle2/topcrt/backend-debug
- AOB?
8-APR-2013
--
iTOP
meeting
(2-3pm)
- Call into Gemba (Brian/Matt A)
- Mechanical preparations/planning (Marc + ...)
- CRT analysis of Elog 95 data [plot]
(Mike)
- Manpower/schedule
discussion
- AOB?
1-APR-2013
--
iTOP
meeting
(2-3pm)
- Call into Gemba (Brian/Matt A)
- Brief Software meeting summary (Matt B)
- LEPS data analysis (Mike) -- deferred
- Manpower/schedule
discussion
- AOB?
18-MAR-2013
--
iTOP
meeting
(2-3pm)
- Call into Gemba (Matt B)
- Schedule
Update (Brian K)
- LEPS data analysis (Mike)
- Regular Updates:
- IRS3B/electronics Testing Update/plans
(Gary)
- Mechanics updates (Marc)
- Sci-Fi Tracker -- update from Fuji hall (Brian K)
- Calibration laser test plans (Bostjan/Matt
A/Gary)
- Note: (this week!)
iTOP Final Electro-mechanical Packaging Task Force [Workshop]
(March 20-21, ID Lab)
- LEPS (SPring-8) beam test May 9 - 18 [schedule
page] [details
on fill modes]
- AOB?
25-FEB-2013
-- continuing into B2GM
4/5-MAR-2013
iTOP
meeting (2-3pm)
- Software
meeting summary (Matt B)
- LEPS data analysis (Mike)
- There will not be a bPID meeting Wednesday (next week is
B2GM):
- IRS3B/electronics Testing Update
(Gary)
- Mechanics updates (Marc)
- Sci-Fi Tracker -- update from Fuji hall
- Calibration Laser system update
(6-MAR-2013) -- difficulties of initial concept identified
(Bostjan)
- TOP TWiki update/presentation (Bostjan)
- Scheduling update for CRT work [link]
- iTOP Final Electro-mechanical Packaging Task Force [Workshop]
(March 20-21, ID Lab)
- LEPS (SPring-8) beam test May 9 - 18 [schedule
page] [details
on fill modes]
- AOB?
11-FEB-2013
iTOP
meeting (2-3pm)
- Software
meeting summary (Matt B)
- LEPS data analysis (Mike)
- There will not be bPID Wednesday (but next Monday is a
holiday):
- IRS3B/electronics Testing Update
(Gary) [20feb
morning TOP presentation]
- Mechanics [diagram]
(Marc)
- Board
Stack [schedule]
(Matt A)
- Sci-Fi Tracker (Brian K) [link]
- Calibration Laser system update (Bostjan) [link]
- CRT Data Rate/volume Estimates [pdf] [xls]
(Gary)
- Scheduling update for CRT work [link]
- iTOP Final Electro-mechanical Packaging Task Force [Workshop]
(March 20-21, ID Lab)
- AOB?
4-FEB-2013
iTOP
meeting (2-3pm)
- Report on Fuji
Hall CRT progress (Gary)
- Software
meeting summary (Matt B)
- LEPS data analysis (Mike)
- There will be bPID Wednesday meeting:
- IRS3B/electronics Testing Update [draft]
(Gary)
- Mechanics (Marc)
- Board
Stack [schedule]
(Matt A)
- Sci-Fi Tracker (Brian K) [link]
- Calibration Laser system [concept]
(Gary)
- Scheduling update for CRT work [link]
- AOB?
28-JAN-2013
iTOP
meeting (2-3pm)
- Report from bPID/EB/Fuji Hall (Gary)
- DOE CR guidance (OPC funding)
- Work plan this week:
- Rudimentary
readout, including 2x CAMAC crates, CR triggered
data
- Confirmation of Event
T0 (event synchronization of FTSW trigger via
CAMAC TDC)
- Software meeting summary (Matt B)
- Summary on
LEPS data analysis (Mike)
- Next bPID Wednesday meeting deferred until Feb 6, but
local updates anyway:
- IRS3B Testing Update (Gary)
- Mechanics (Marc)
- Board Stack (Matt A)
- Sci-Fi Tracker (Brian K)
- Calibration Laser system (Gary)
- Other information?
- Continued Scheduling discussion for CRT work (travel
dates fixed?)
- AOB?
18-JAN-2013
Firmware meeting with
SLAC (11am)
- CTA-related items
- Q&A regarding trigger studies in
Hawaii
- bypassing/VR improvements
- Belle II related items (Kurtis and Luca will join)
- IRS3B firmware (brief) status [Luca]
- discussion on testing/task sharing
- brief fDIRC firmware status update
- AOB (?)
- Current firmware
repository for the IRS2 version of the board stack
http://idlab-scrod.googlecode.com
svn path: svn/SCROD-boardstack/new_daq_interface
*some instructions still coming to deal with a couple compile
"features"
**note that this
includes schematics for various boards here:
https://code.google.com/p/idlab-scrod/downloads/list
- Data
formats for this version of the firmware:
http://www.phys.hawaii.edu/~kurtisn/doku.php?id=itop:documentation:data_format
- Schematics
for the IRS3B version of the carrier levels 0 and 2
http://www.phys.hawaii.edu/~mza/PCB/iTOP/carriers/carrier02.revB.schematic.pdf
14-JAN-2013
iTOP
meeting (2-3pm)
- Intro -- Bostjan Macek (will be in?)
- Software meeting summary (Matt B)
- Updates from Japan (Tom)
- Updates on LEPS data analysis (Mike)
- Updates for this week's morning bPID:
- IRS3B Testing Update (Gary)
- Mechanics (Marc)
- Board Stack (Matt A)
- Sci-Fi Tracker (Brian K)
- Other information?
- Preparations for the TrgDAQ meeting next week (Brian will
run meeting?)
- Continued Scheduling discussion for CRT work
- AOB?
7-JAN-2013
iTOP
meeting (2-3pm)
- Software meeting summary (Matt B)
- Updates on LEPS data analysis (Mike)
- Updates for this week's morning bPID:
- IRS3B Testing [Draft]
(Gary)
- Mechanics (Marc)
- Summary of Board Stack status/schedule - status
of carrier boards (Matt A)
- Sci-Fi Tracker schedule update (Brian K)
- LEPs data analysis (Mike)
- Continued Scheduling discussion for CRT work (pushing
into February)
- AOB?
10-DEC-2012
iTOP
meeting (2-3pm)
Summary
- Software meeting summary (Matt B)
- Updates:
- Mechanics (Marc)
- Summary of Fuji Hall work (Matt A)
- Sci-Fi Tracker schedule update (Brian K)
- LEPs data analysis (Mike)
- Continued Scheduling discussion for autumn/winter CRT
work (pushing into January)
- AOB?
3-DEC-2012
iTOP
meeting (2-3pm)
- Software
meeting summary (Matt B)
- Updates:
- Mechanics (Marc)
- Thermal/mechanical housing/mock-up plans &
schedule
- Laser calibration system design specifications
(simple diffuser/quartz block test?)
- Comment on (inverted) assembly/lift
stage
- Readout/news from Fuji Hall (Kurtis/Matt A)
- Analysis of LEPS
data Exp. 1&2 and description
(Mike)
- Carrier update: [schematics]
and comments thereon
as well as mentioned .ucf
(Matt A)
- HV (first
module received), as well as comments on use
and mechanical notes
(Gerard)
- Sci-Fi Tracker schedule update (Brian K)
- Continued Scheduling discussion for autumn/winter CRT
work (pushing into January)
- Matt A 11/18-12/9; Kurtis 11/17-12/7;
- Plans for testing with what will be available.
- AOB?
26-NOV-2012
iTOP
meeting (2-3pm)
- Software
meeting summary (Matt B)
- Updates from KEK:
- Mechanics (Marc)
- Comments
on Jim Fast's horseshoe/module concept
- Wedge mechanical dimensions
- Thermal/mechanical housing/mock-up plans &
schedule
- Laser calibration system design specifications
(simple diffuser/quartz block test?)
- Readout (Kurtis/Matt A)
- Regular updates (what to present @ morning PID) ?
- Analysis of LEPS data and description (Mike)
- Carrier update? (Matt A)
- IRS3B Testing (Gary)
- Sci-Fi Tracker schedule update (Brian
K)
- Continued Scheduling discussion for autumn/winter CRT
work (pushing into January)
- Matt A 11/18-12/9; Kurtis 11/17-12/7;
- Plans for testing with what will be available.
- AOB?
19-NOV-2012
iTOP meeting (2-3pm)
- Updates from KEK
- Mechanics
(Marc)
- CRT Slider mech reference
- Readout (Kurtis/Matt A)
- Regular updates
- Analysis of LEPS data and description (Mike)
- Carrier update? (Matt A)
- Sci-Fi Tracker schedule
update (Brian K)
- Continued Scheduling discussion for autumn/winter CRT
work (pushing into January)
- Matt A 11/18-12/9; Kurtis
11/17-12/7; Marc 11/15-11/21
- Plans for testing with what will be available.
- AOB?
05-NOV-2012
iTOP
meeting (2-3pm)
Draft
B2GM TOP Electronics presentation
- Software
meeting summary (Matt B)
- Regular updates
- Analysis of LEPS data (Mike)
- Scheduling
update on Cosmic Ray Test stand at Fuji Hall (revised
29-oct-2012)
- Carrier update?
- HV testing underway at Indiana
- Sci-Fi Tracker schedule update
- Requests/coordination for iTOP B2GM talks
- Electronics update - Gary
- General CRT update?
- Mechanical?
- Scheduling?
- Continued Scheduling discussion for autumn/winter CRT
work
- Dedicated mechanical/working meeting in Fuji hall
--> "in conjunction with B2GM"
- Gary: 11/2 - 11/18; Matt A 11/18-12/9;
Kurtis
11/17-12/7; Matt B. 11/10-11/23; Tom 11/11-11/18;
Marc 11/15-11/21
- Items for Matt B. or Tom bring for Gary
- Plans for testing with what will be available.
- Others to be decided: Brian K(?)
- AOB?
29-OCT-2012
iTOP
meeting (2-3pm)
- Software
meeting summary (Matt B)
- US Hardware/Firmware meeting summary (Brian)
- Regular updates -- in preparation for Wednesday meeting
- Analysis of LEPS data (Mike)
- Scheduling
update on Cosmic Ray Test stand at Fuji Hall (revised
29-oct-2012)
- Fuji Hall Equipment
list
draft (24-OCT-2012) [Kurtis/Matt]
- IRS3B Testing Update (Gary) [draft
slides]
- Strongback
analysis -- response from Nagoya? (Marc)
- Mechanical/thermal Mock-up set-up in WAT 221
(all)
- New Carrier Boards and Front-boards update
(Matt A)
- HV update (Gerard) [repository]
- 5 day turn (available Nov. 2 or 5)
- Quote for 70 boards (production run) only slightly
more than prototype run
- Likes RG-316, but wants to do a burn-in test
- Sci-Fi Tracker [schedule update]
- Firmware: TARGET4/readout
(Xiaowen/Brian)
- KLM MB status (Xiaowen)
- Support mechanics/assy & test plans (all)
- Plans for TARGET5 & TARGET6
- Continued Scheduling discussion for autumn/winter CRT
work
- Dedicated mechanical/working meeting in Fuji hall
--> "in conjunction with B2GM"
- Gary: Nov. 2 - 18; Matt A Nov.
11/10-30; Kurtis 11/16-12/7;
- Plans for testing with what will be available.
- Others to be decided: Marc, Brian K(?), Matt
B(?), Tom(?)
- AOB?
22-OCT-2012
iTOP
meeting (2-3pm)
- Software
meeting summary (Matt B)
- Draft Firmware Self-Review (Kurtis)
- Regular updates -- in preparation for Wednesday
meeting: [electronics
draft]
- Analysis of LEPS data (Mike)
- Scheduling
update on Cosmic Ray Test stand at Fuji Hall (revised
23-oct-2012)
- Fuji Hall Equipment
list draft (24-OCT-2012) [Kurtis/Matt]
- IRS3B Testing Update (Gary)
- Strongback
analysis -- response from Nagoya? (Marc)
- Mechanical Mock-up set-up in WAT 221 (all)
- New Carrier Boards and Front update (Matt A)
- Thermal mitigation:
- Coupling heat into side walls (Matt A/Gary)
- Integrated heat-pipes/manifold (Marc)
including mock-up
- Sci-Fi Tracker
- Firmware: TARGET4/readout
(Xiaowen/Brian)
- KLM MB status (Xiaowen)
- Support mechanics/assy & test plans (all)
- Plans for TARGET5 & TARGET6
- Continued Scheduling discussion for autumn/winter CRT
work
- Dedicated mechanical/working meeting in Fuji hall
--> "in conjunction with B2GM"
- Gary: Nov. 2 - 18; Matt A Nov.
11/10-30; Kurtis 11/16-12/7;
- Plans for testing with what will be available.
- Others to be decided: Marc, Brian K(?), Matt
B(?), Tom(?)
- AOB?
15-OCT-2012
iTOP
meeting (2-3pm)
- Software meeting summary (Matt B)
- Scheduling
update (Kurtis/Gary/Matt A)
- LEPS run redux (Matt B)
- Regular updates -- in preparation for Wednesday
meeting: [electronics
draft]
- Analysis of LEPS data (Mike)
- Update on Cosmic Ray Test stand at Fuji Hall
- Schedule update (Gary)
- Strongback analysis -- response from Nagoya?
(Marc)
- Figure from
Suzuki-san [Kohriki-san] on joint QBB/SB
(14-OCT-2012)
- Mechanical Mock-up set-up in WAT 221 (all)
- Board-stack update:
- Firmware review report(?) (Kurtis)
- New Carrier Boards update (Matt A)
- Amplifier simulations [OP846/OP847 options]
(Gary/Luca)
- Front-front/front-back --> "Front" board
(Matt A)
- Gerard on HV schedule [xls]
- Belle2Link meeting
- Thermal mitigation (proposed
partition) latest power update [xls]
- Coupling heat into side walls (Matt A/Gary)
- Integrated heat-pipes/manifold (Marc)
including mock-up
- Sci-Fi Tracker
- Firmware: TARGET4/readout
(Xiaowen/Brian)
- KLM MB status (Xiaowen)
- HV Rack set-up in WAT 221 (all)
- Support mechanics/assy & test plans (all)
- Report from TARGET6 Review (Gary)
- IRS3B Testing Status (Gary)
- Continued Scheduling discussion for autumn/winter CRT
work
- Dedicated mechanical/working meeting in Fuji hall
--> "in conjunction with B2GM" (will mechanics be
ready??)
- Gary: Nov. 2 - 18; Matt A Nov.
11/10-30;
Kurtis 11/16-12/7;
- Will CRT mechanics be ready??
- Others to be decided: Marc, Brian K(?), Matt
B(?), Tom(?)
- AOB?
8-OCT-2012
iTOP
meeting (2-3pm)
- Feedback from Special BPAC Review (Oct. 1-2 @ KEK)
(Gary)
- Software
meeting summary? (Matt B away at LEPS, though
meeting to be held this morning)
- LEPS run status (Matt B)
- Regular updates:
- Updates on
FNAL studies [txt]
(Mike)
- Update on Cosmic Ray Test stand at Fuji Hall
- Strongback analysis [pdf]
(Marc)
- Mechanical Mock-up set-up in WAT 221 (all)
- Board-stack update:
- Firmware (Kurtis)
- New Carrier Boards (Matt A)
- Amplifier simulations [png]
(Luca)
- Front-front/front-back? (all)
- Any news from Gerard on HV? (all)
- Belle2Link ping (Gary)
- Thermal mitigation (proposed
partition)
- Coupling heat into side walls (Matt
A/Gary) including modelling [pdf]
- Integrated heat-pipes/manifold (Marc)
including mock-up
- Sci-Fi Tracker
- Firmware: TARGET4/readout
(Xiaowen/Brian)
- KLM MB status (Xiaowen)
- HV Rack set-up in WAT 221 (all)
- Support mechanics/assy & test plans (all)
- IRS3B Testing Status (Gary)
- Continued Scheduling discussion for autumn/winter CRT
work
- Dedicated mechanical/working meeting in Fuji hall
--> "in conjunction with B2GM"
- Working schedule/costs with Indiana for HV part of
board stack
- Gary: Nov. 2 - 18; Matt A Nov.
11/10-30;
Kurtis 11/16-12/7;
- Others to be decided: Marc, Brian K(?), Matt
B(?), Tom(?)
- TARGET6 Design Review, Belle II Firmware @ SLAC
Friday, Oct. 12, 2012
- AOB?
24-SEP-2012
iTOP
meeting (2-3pm)
- Brief software meeting summary (Matt B)
- Regular updates:
- Updates on
studies (Mike)
- Update on Cosmic Running at Fuji Hall
- Optimization of Test stand/detector configuration
(Matt B)
- Strongback analysis, other mechanical preparation
(Marc)
- Sci-Fi Tracker
- Support mechanics (Gary/Marc)
- KLM MB status (Xiaowen)
- Firmware: TARGET4/readout
(Xiaowen/Brian)
- IRS3B (Gary)
- IRS3B board-stack daughtercard
- Universal Eval Rev. B daughtercard
- US Hardware/Firmware meeting this Friday
- Oct. 2-9 LEPS run status?
- Continued Scheduling discussion for autumn/winter CRT
work
- Dedicated mechanical/working meeting in Fuji hall
--> "in conjunction with B2GM"
- Working schedule/costs with Indiana for HV part of
board stack
- Gary: Nov. 2 - 18; Matt A Nov.
11/10-30;
Kurtis 11/16-12/7;
- Others to be decided: Marc, Brian K(?), Matt
B(?), Tom(?)
- Preparations for upcoming Reviews:
- Belle II Focused Review (PXD, PID) Oct. 1-2 @ KEK
- Draft slides posted on ASIC/readout
- Manpower meeting at KEK 9/30
- AOB?
17-SEP-2012
iTOP
meeting (3-4pm)
- Brief software meeting summary (Matt B)
- Regular updates:
- Updates on
studies (Mike)
- Update on Cosmic Running at Fuji Hall
- Optimization of Test stand/detector configuration
(Matt B)
- Strongback analysis, other mechanical preparation
(Marc)
- Sci-Fi Tracker
- Comments on
previous acquisition/mechanics (Marc)
- Mechanics/photonics (Gary)
- KLM MB status (Xiaowen)
- Firmware: TARGET4/readout
(Xiaowen/Brian)
- IRS3B (Gary)
- IRS3B board-stack daughtercard
- Universal Eval Rev. B daughtercard
- Oct. 2-9 LEPS run status?
- Continued Scheduling discussion for autumn/winter CRT
work
- Dedicated mechanical/working meeting in Fuji hall
--> "in conjunction with B2GM?"
- Working schedule/costs with Indiana for HV part of
board stack
- Matt A away a bit longer...
- Gary: Nov. 2 - 18; Matt A Nov.
11/10-30;
Kurtis 11/16-12/7;
- Others to be decided: Marc, Brian K(?), Matt
B(?), Tom(?)
- Preparations for upcoming Reviews:
- Belle II Focused Review (PXD, PID) Oct. 1-2 @ KEK
- Really no other reviews this calendar year??
- AOB?
10-SEP-2012
iTOP
meeting (3-4pm)
- Software meeting summary (Matt B)
- bPID: preparations for this
week's mechanical meeting (Tuesday meeting slot?)
- Regular updates:
- Updates on
studies -- deferred as Mike is out (Mike)
- Update on Cosmic Running at Fuji Hall
- Optimization of Test stand/detector configuration
(Matt B)
- Strongback analysis, comments
[in advance of tomorrow's meeting] (Marc)
- NB: still awaiting feedback from Nagoya on
how to fasten strongback to the QBB
- Oct. 2-9 LEPS run status?
- Continued Scheduling discussion for autumn/winter CRT
work
- Dedicated mechanical/working meeting in Fuji hall
--> "in conjunction with B2GM?"
- Working schedule/costs with Indiana for HV part of
board stack
- Matt A away a bit longer...
- Gary: Nov. 2 - 18; Matt A Nov.
11/10-30;
Kurtis 11/16-12/7;
- Others to be decided: Marc, Brian K(?), Matt
B(?), Tom(?)
- Preparations for upcoming Reviews:
- Belle II Focused Review (PXD, PID) Oct. 1-2 @ KEK
- Really no other reviews this calendar year??
- AOB?
20-AUG-2012
iTOP
meeting (3-4pm)
- Software
meeting summary (Matt B)
- bPID: preparations for this week's meeting
- Updates on
studies (Mike)
- Update on Cosmic Running at Fuji Hall
- Optimization of Test stand/detector configuration
(Matt B)
- Structural analysis (Marc)
- fDIRC [firmware] development
schedule/status/plans (Kurtis)
- Readout module/hardware schedule/status/plans
(Matt A)
- Oct. 2-9 LEPS run status?
- Continued Scheduling discussion for autumn/winter CRT
work
- Dedicated mechanical/working meeting in Fuji hall
--> "after LEPS run"
- Gary: Nov. 2 - 18; Matt A Nov.
11/10-30;
Kurtis 11/16-12/7;
- Others to be decided: Marc, Brian K(?), Matt
B(?), Tom(?)
- SLAC Beam Test option
(request) (Gary -- Kurtis to present on
Wednesday?)
- Preparations for upcoming Reviews:
- ESTB
First
meeting -- 23-AUG-2012 (this Thursday)
- ANL
System
Review -- 24-AUG-2012 (this Friday!!)
- Belle II Focused Review (PXD, PID) Oct. 1-2 @ KEK
- AOB?
13-AUG-2012
iTOP
meeting (3-4pm)
- Software
meeting summary (Matt B)
- bPID: no meeting this week
- Updates on
studies (Mike)
- Mechanics/readout/commissioning for Cosmic
Running at Fuji Hall
- Participation in Oct. 2-9 LEPS run (Tom?
Matt B)
- Schedules for autumn CRT work
- Dedicated mechanical/working meeting in Fuji hall
(end of August/early Sept) --> "after LEPS run"
- Gary: Nov. 2 - 18; Matt A Nov.
11/10-30;
Kurtis 11/16-12/7; Brian K(?), Matt B(?)
- Preparations for upcoming Reviews:
- ANL System Review
-- 24-AUG-2012 (next Friday!!)
- Belle II Focused Review (PXD, PID) Oct. 1-2 @ KEK
- Updates?
- Update on Rev. A2 board-stack for CRT (Matt A)
- Update on firmware development plans [fDIRC] (Matt
A/Gary)
- Plans for declaring T-1019 data analysis ?? (news
from PNNL?)
- AOB?
6-AUG-2012
iTOP
meeting (3-4pm)
- Software
meeting summary (Matt B)
- bPID meeting this week -- material to present
- Updates on
studies (Mike)
- Mechanics for Cosmic Running at Fuji Hall
- Constraints from simulation (Matt B)
- Comments
on posted mechanics materials (Marc)
- Dedicated Mechanics Meeting
- Follow-up?? (call for separate meeting during
this weeks one??)
- Participation in Oct. 2-9 LEPS run (Matt B)
[next time]
- Schedules for autumn CRT work
- Dedicated mechanical/working meeting in Fuji hall
(end of August/early Sept) --> "after LEPS run"
- Gary: Nov. 2 - 18; Matt A Nov.
11/10-30;
Kurtis 11/16-12/7; Brian K(?), Matt B(?)
- Preparations for upcoming Reviews:
- ANL System Review (dates TBD -- doodle poll, awaiting
ANL confirmation)
- Belle II Focused Review (PXD, PID) Oct. 1-2 @ KEK
- Deferred?
- Update on Rev. A2 board-stack for CRT (Matt A)
- Update on firmware development plans [fDIRC] (Kurtis)
- Plans for T-1019 data reprocessing ?? (news from
PNNL?)
- AOB?
30-JUL-2012
iTOP
meeting (3-4pm)
- Next Software meeting is 2-3pm Monday, August 6th (Matt
B)
- Feedback from Trg/DAQ, B2GM (Gary)
- Commissioning schedule discussion [PDF]
- Technical issues [PDF]
- Schedule [PDF]
- bPID meeting this week -- material to present
- Updates on
studies (Mike)
- Mechanics for Cosmic Running at Fuji Hall
- Constraints from simulation (Matt B) [next
time]
- Division of labor (PNNL/UH) on CRT mechanics
(discussion)
- Dedicated Mechanics Meeting
- Status of Rev. A2 board-stack for CRT (Matt
A)
- Participation in Oct. 2-9 LEPS run (Matt B)
[next time]
- Schedules for autumn CRT work
- Dedicated mechanical/working meeting in Fuji hall
(end of August/early Sept)?
- Gary: Nov. 2 - 18; Matt A Nov.
11/10-30;
Kurtis 11/16-12/7; Brian K(?), Matt B(?)
- Preparations for upcoming Reviews:
- ANL System Review (dates TBD -- doodle poll)
- Belle II Focused Review (PXD, PID) Oct. 1-2 @ KEK
- Deferred:
- Update on firmware development plans [fDIRC] (Kurtis)
- Plans for T-1019 data reprocessing ??
- AOB?
25-JUN-2012
iTOP
meeting (3-4pm)
- 2-3 pm is Software
meeting (Matt B)
- Revised CD-1 Review date ("tomorrow!!" Tuesday,
June 26, 2012)
- Status and preparation
- Material into breakout slides (for discussion -- 2.5
hours for "everything")
- bPID meeting this week -- material to present
- Updates on
studies (Mike)
- Mechanics for Cosmic Running at Fuji Hall
(Marc)
- Dedicated Mechanics Meeting?
- Latest QBB mechanical design from Kohriki-san [link]
- Where the "20120619" version is for prototypes to
be assembled this summer and are shown in 2D CAD.
- The "20120622" version is thecurrent best QBB
design considered for the real counter module and shown
in 3D CAD.L
- Fast BB module [pdf]
- Summary of discussion from last week
- Update on bench test measurements (Kurtis)
- Reprocessed data analysis for T-1019; strategy &
status?? (all)
- IRS3B Design Review (15-JUN-2012) and Submission
(18-JUN-2012); Tape-out
(25-JUN-2012)
[link]
- Summary of Review [link]
- Discussion of plans
- Evaluating external timebase PLL/DLL circuit
- Prep IRS3B eval board prior to arrival of IRS3B
- Improvements to HV power: possible
active circuit for SL-10 (Gerard -- 21-JUN-2012) [PDF]
- And some follow-on questions
- Plans for upgraded electronics/commissioning @ KEK:
- in "summer" -- 1x full module of IRS2 w/ AD8000
amps [fDIRC deployment as test bed]
- in "autumn" -- 1x full module of IRS3B w/ AD8000
amps [IRS3B ready "mid September"]
- (TARGET5 in fab, due 30-AUG-2012; TARGET6 end
of this year)
- TARGET4-based SciFi tracker; 9U VME board being
designed
- Schedule for commissioning/integration
- Preparations for upcoming B2GM Meetings
- AOB?
21-MAY-2012
iTOP
meeting (3-4pm)

- 2-3 pm is Software meeting (Matt B)
- Revised date for CD-1 ("late June" 2012)
- Practice call tomorrow (management predominantly...)
- Material into breakout slides (for discussion -- 2.5
hours for "everything")
- bPID meeting this week -- material to present
- Updates on
studies (Mike)
- Mechanics for Cosmic Running at Fuji Hall
(Marc)
- Tent diagram from Hayakawa-san for discussion [pdf]
- Dedicated mechanical meeting?
- Update on bench test measurements (Kurtis)
- Reprocessed data analysis for T-1019; end of May
deadline? (all)
- Plans for upgraded electronics/commissioning @ KEK:
- in "summer" -- 1x full module of IRS2 w/ AD8000 amps
- in "autumn" -- 1x full module of IRS3B w/ AD8000 amps
- IRS3B design review 15-JUN @ SLAC [design
changes fixed Friday]
- (TARGET5 to be fabbed in 4-JUN run; TARGET6 end
of this year)
- TARGET4-based SciFi tracker
- Schedule for commissioning/integration
- Future:
- Another beam test...
- AOB?
7-MAY-2012
iTOP
meeting (3-4pm)
- 2-3 pm is Software meeting (Matt B) [Gary @
Mike Cooney Oral exam]
- Items to prepare for upcoming CD-1 (31-MAY-2012) -- WAS DUE
FRIDAY (as is DOE Renewal...)
- Revised CDR (incorporate appendix material, one new
measurement)
- Material into breakout slides (for discussion -- 2.5
hours for "everything")
- bPID meeting this week -- material to present
- Updates on
studies (Mike)
- Mechanics for Cosmic Running at Fuji Hall
(Marc)
- Update on bench test measurements (Kurtis)
- Reprocessed data analysis for T-1019; end of May
deadline (all)
- Plans for upgraded electronics/commissioning @ KEK:
- in "summer" -- 1x full module of IRS2 w/ AD8000 amps
- in "autumn" -- 1x full module of IRS3B w/ AD8000 amps
- Future:
- Cosmic Ray test schedule
- Another beam test...
- AOB?
30-APR-2012
iTOP
meeting (3-4pm)
- 2-3 pm is Software meeting (Matt B)
- Mechanical Engineering/Integrated Module preparation for
upcoming CD-1 (31-MAY-2012)
- SCROD Rev B/board-stack 3D mechanics (Matt A)
- Board stack for breakout?
- No bPID meeting this week (Nagoya @ LEPS/Golden Week)
- Nphoton studies [30-APR-2012
edition] (Mike)
- Mechanics for Cosmic Running at Fuji Hall
(Marc)
- Update on bench test measurements (Kurtis)
- Reprocessed data analysis for T-1019; end of May
deadline (all)
- Further discussion on schedule for upgraded
electronics/commissioning @ KEK in "summer"
- Future:
- Cosmic Ray test schedule
- Another beam test...
- AOB?
23-APR-2012
iTOP
meeting (3-4pm)
- 2-3 pm is Software
meeting (Matt B)
- Mechanical Engineering/Integrated Module preparation for
upcoming CD-1 (mid-May)
- SCROD Rev B/board-stack 3D mechanics (Matt A)
- Thermal modelling plans ?
- No bPID meeting this week (Nagoya @ LEPS) -- local
updates
- Summary of studies (Mike)
- Mechanics for Cosmic Running at Fuji Hall
(Marc)
- Update on bench test measurements (Kurtis)
- Schedule for upgraded electronics/commissioning @ KEK
in "summer"
- CDR updates
(Gary/Tom/Kurtis)
- Future:
- Major push on beam test data analysis/additional cosmic
running
- Cosmic Ray test schedule
- Another beam test...
- AOB?
9-APR-2012
iTOP
meeting (3-4pm) -- note new time
- 2-3 pm is Software
meeting (Matt B)
- Mechanical Engineering/Integrated Module preparation for
upcoming CD-1 (mid-May)
- SCROD Rev B/board-stack 3D mechanics (Matt A)
- Thermal modelling plans
- Items for bPID meeting this week
- Summary of studies (Mike) [txt]
- Mechanics for Cosmic Running at Fuji
Hall (Marc)
- Update on bench test measurements (Kurtis)
- Schedule for upgraded electronics/commissioning @ KEK
in "summer"
- CDR
updates (Gary/Tom/Kurtis)
- Future:
- Major push on beam test data analysis/additional cosmic
running
- Another beam test...
- Review schedule together
- AOB?
5-MAR-2012
iTOP
meeting (2-3pm)
- Mechanical Engineering/Integrated Module preparation for
upcoming reviews (Tom/Marc/Matt)
- Concept drawings of electronics?
- Thermal modelling?
- Items for bPID meeting this week (ongoing T-1019
analysis)
- Laser studies (Mike) [PDF]
- Update on Nphoton (Gary) [E-log
link] : #43
- Update on bench test measurements (Kurtis)
- Software
meeting after this (Matt B)
- CDR
updates (Gary/Tom/Kurtis)
- Future:
- Major push on beam test data analysis/additional cosmic
running
- Renewed Cosmic
Running at Fuji Hall
- Another beam test...
- Review schedule together
- AOB?
- Deferred: grounding discussion based upon summary
6-FEB-2012
iTOP
meeting (2-3pm)
- Mechanical Engineering preparation for upcoming reviews
(Tom/Marc)
- Mechanical mock-up plans (Matt, Kurtis away...)
- Concept drawings of support/housing sealing (Jim
F.) [prism1] [prism_detail]
- FEA analysis
- Items for bPID meeting this week?
- Specifications for mirrors/wedges
- Update on beam test analysis...
- Further Observations on the T-1019 data [E-log
link] (Mike)
- Software
meeting after this (Matt B)
- WBS updates
PNNL (Gary)
- Future:
- Major push on beam test data analysis/additional cosmic
running
- Another beam test...
- Review schedule together
- AOB?
30-JAN-2012
iTOP
meeting (2-3pm)
- Mechanical Engineering preparation for upcoming reviews
(Tom/Marc)
- Mechanical mock-up?
- Need concept drawing of support/housing sealing
- FEA analysis to be performed by Metal
Technology
Inc. of [these
structures?]
- Observations on the T-1019 data (Mike)
- Software meeting tomorrow (Matt B)
- Specs
table to work -- and expanded list [link]
- Are all topics
covered?
- Do we have a schedule that gets us through various
reviews?
- Items for bPID meeting this week?
- Specifications for mirrors/wedges
- Photon losses at window [staric example]
- closeout frame versus window [Ushiroda example]
- Step-facet the window? [Alan
sketch]
- Update on beam test analysis...
- Feedback from
WBS Meetings at PNNL (Gary)
- WBS resource leveling
- Updated WBS (substantially) and Risk Assessment
- Critical issue -- understanding timing peformance
- Ready for end of March Director's Review "do over"?
- AOB?
- Deferred:
- Another beam test...
- Review schedule together
9-JAN-2011
iTOP
meeting (2-3pm)
- Laser anomolies in the T-1019 data (Mike)
- Software meeting tomorrow? (Matt B away)
- Belle II Trigger/DAQ
meeting here [Jan. 16-18]
- Plans for
Director's Review (anon)
- WBS resource levelling
- Visit to
- Another beam test...
- Next bPID
meeting?
28-NOV-2011
iTOP
meeting (2-3pm)
- Latest updates in the Beamtest/MC comparison saga (Mike)
- Laser degradation during run? [txt]
- Comparison plots [pdf]
- BPAC summary notes [txt] (Gary)
- Software meeting (Matt B)
- T-1019 preparations (Gary)
- Update regarding Nagoya status/shipment
[Kurtis/Gary]
- SciFi tracker update (Dmitri here) status/work plan
- A number of update items
- ORC
due tomorrow
- Critical items list/packing
- Plans for
Director's Review (1/3 technical, 2/3 budget & schedule)
- Quotation preparation
- Slides due Friday
21-NOV-2011
iTOP
meeting (2-3pm)
- Latest updates in the Beamtest/MC comparison saga [pdf] (Mike)
- Software meeting deferred -- due to meeting at KEK
(Matt B)
- T-1019 preparations (Gary)
- Update from Nagoya [Gary/Kurtis]
- SciFi tracker status/work plan
- A number of update items
- PREP request & other follow-up
- Plans for
Director's Review (1/3 technical, 2/3 budget & schedule)
14-NOV-2011
iTOP
meeting (2-3pm)
No
meeting this week due to BPAC, BGM, GBC, B2GM
7-NOV-2011
iTOP
meeting (2-3pm)
- Latest updates in the Beamtest/MC comparison saga [run 310 timing] (Mike)
- Software/BPAC figures update [ repository
] (Matt B)
- T-1019 preparations (Gary)
- Update from Nagoya [Kurtis?] (Matt A
arrived back this morning)
- SciFi tracker mounting [fig]
- 13 or so Account renewal Approvals issued
- PREP request
- BPAC practice talks @ 2pm tomorrow [review
link]
- Electronics and trigger status 15'
[11:50]
Speaker: |
G. Varner (Univ. of Hawaii) |
- Calibration (T0 and laser
calibration) 15' [14:10]
Speaker: |
K. Nishimura (Univ. of Hawaii) |
- Quartz procurement 15'
[15:00]
- Various travel upcoming... (all)
- WBS 2.0
(integration/commissioning) update needed (in preparation
Director's Review)
- AOB?
31-OCT-2011
iTOP
Halloween meeting (2-3pm)
- Latest updates in the Beamtest/MC comparison saga
(Mike)
- Software meeting @ ???? tomorrow (Matt
B) [link]
- T-1019 preparations (Gary)
- Mechanics of pivoting stage
- What survey equipment do we have/need?
- bPID Meeting @ 2pm tomorrow [link]
- The agenda is as follows:
- Update on 2010 beam test analysis M.Jones, Y.Arita, K.Inami
- Simulation studies for cosmic-ray/beam test M.Barrett
- Cosmic-ray test stand status K.Suzuki, Y.Horii, Y.Koga,
N.Kiribe
- MCP-PMT integration status K.Matsuoka, H.Takeichi, S.Hirose
- Electronics updates M.Andrew, K.Nishimura
- Discussion for BPAC, beam test
- Nagoya integration update and plans? (Matt A in
Nagoya)
- SciFi tracker opto/electro/mechanics updates (Marc)
- BPAC preparations (all) [review
link]
- Various travel upcoming... (all)
- WBS 2.0
(integration/commissioning) update needed (in preparation
Director's Review)
- AOB?
24-OCT-2011
iTOP
meeting (2-3pm)
- Latest updates in the Beamtest/MC comparison saga
(Mike)
- Software meeting @ 3pm tomorrow (Matt B) [link]
- Nagoya integration update and plans? (Matt A en
route Wednesday to Nagoya)
- Access window update (Inami) [pdf]
- SciFi tracker opto/electro/mechanics updates (Marc/Gary)
- BPAC preparations (all) [review
link]
- Various travel upcoming... (all)
- T-1019 preparations (Gary)
- Mechanics of pivoting stage (total weight
estimate?)
- Support for trigger counters (query to Inami)
- What survey equipment do we have?
- An issue from the quartz RFP (qualifying the optics)
- WBS 1.3
(electronics) update; WBS 2.0
(integration/commissioning) update needed
- AOB?
17-OCT-2011
iTOP
meeting (2-3pm)
- Latest update in the Beamtest/MC comparison saga
(Mike)
- Channel numbering issue? [txt]
- Slides [pdf]
- Software meeting @ 3pm tomorrow (Matt B)
[link]
- Nagoya integration update?
- Items for Wednesday's bPID meeting?
- Update on 2010 beam test MC/data comparison
(Mike)
- US/CosRay simulation update (Matt B)
- Electronics update (Matt A?)
- T-1019 update (Gary)
- Response due by Oct 25 to Access window questions for
ODHM:
- Suzuki-san's message [txt]
- Link to last ODHM [link]
- Kohriki-san's drawing [pdf]
- IDS backward view [pdf]
- SciFi tracker opto/electro/mechanics update (Gary)
- BPAC preparations (all) [review
link]
- T-1019 updates (Gary)
- updated MOU (to go out for signatures this week)
[pdf]
- T-1019 homepage [link]
- Estimate of optics assembly effort (v0.2) [docx] [pdf]
- AOB?
10-OCT-2011
iTOP
meeting (2-3pm)
- Latest in the Beamtest/MC comparison saga [72.5 degree
incident angle] (Mike)
- Description text [txt]
- MC-data comparison Plots [pdf]
- Software meeting @ 3pm tomorrow (Matt B) [link]
- Nagoya integration (Kurtis/Matt/Marc)
- Kurtis report [pdf]
- Action items (elec):
- Interconnect Rev. B (review this afternoon)
- Front-front HV Rev. C arcing (review this
afternoon w/ Nagoya, PNNL?)
- Action items (mech)?
- Next round integration deployments
- SciFi tracker electro/mechanics, schedule (Gary/Marc)
- T-1019 updates (Gary)
- AOB?
3-OCT-2011
iTOP
meeting (2-3pm)
- Previous Beamtest/MC comparison update (Mike)
- Description text [txt]
- MC-data comparison Plots [pdf]
- Software meeting @ 3pm tomorrow (Matt B) [link]
- Couldn't make it today -- but link for tomorrow
provided
- root file from Matsuoaka-san providing measured QEs for
PMTs, to input into simulation
- Matched values to be used in MC once tube positions
known (make request to fix during Wednesday's meeting?)
- Nagoya integration push de-brief (Marc/Matt A)
- Integration mechanics (Marc) [pdf]
- First go at Thermal Management (Marc) [pdf]
- Update from Nagoya [pdf]
- SciFi tracker electro/mechanics, schedule (Gary/Marc)
- Integration -- one more time ...--> YES.
Travel requests sent.
- What to present on Wednesday?
- Mechanics update (Marc)
- Electronics update (Matt A/Kurtis)
- US software development (Matt B)
- T-1019 preparations (Gary)
- Beam
test
preparations
- AOB?
12-SEP-2011
iTOP
meeting (2-3pm)
- Software meeting @ 3pm tomorrow
- Mechanics for readout, expansion block support and
mounting (Marc)
- SciFi tracker electro/mechanics (Gary/Marc)
- Integration schedule discussion (all)
- Gary (9/16 - 9/26) [Lynn Wood (9/18 - 9/25)]
- Kurtis (9/21 - 10/7) [Note: @ BAM 9/25 -
9/30]
- Matt A (9/23 - 10/2)
- Marc (9/23 - 10/2)
- AOB?
[NOTE
unusual
date]
6-SEP-2011
iTOP
meeting (2-3pm)
- Software meeting @ 3pm today
- Usual suspects for Thursday 3:30pm iTOP meeting with
Nagoya? (Gary)
- Analysis update (Mike)
- Mechanical update (Marc)
- Electronics update/schedule (Kurtis/Matt A)
- Simulation for CosRay test (Matt B)
- Anything else? (e.g. SciFi update(?) Gary)
- Mechanics for readout, expansion block support and
mounting (Marc)
- SciFi tracker electro/mechanics (Gary/Marc)
- Integration schedule discussion (all)
- Gary (9/16 - 9/26) [Lynn Wood (9/18 - 9/25)]
- Kurtis (9/21 - 10/7) [Note: @ BAM 9/25 -
9/30]
- Matt A (9/23 - 10/2)
- Marc (9/23 - 10/2)
- US firmware meeting w/PNNL this Friday (depart 1 week
thereafter)
- AOB?
29-AUG-2011
iTOP
meeting
- Software meeting tomorrow (prep?)
- Mechanics for readout, expansion block support and
mounting (Marc)
- SciFi tracker electro/mechanics (Gary/Marc)
- Overall design
- Schematics
- Layout
- Integration schedule discussion (all)
- Gary (9/16 - 9/26) [Lynn Wood (9/18 - 9/25)]
- Kurtis (9/21 - 10/7) [Note: @ BAM 9/25 -
9/30]
- Matt A (9/23 - 10/2)
- Marc (when mechanics will be ready?) --
NOTE: need 45 days advanced notice (in principle)
- US hw/fw meeting this Friday (Gary)
- AOB?
22-AUG-2011
iTOP
meeting
- Software meeting tomorrow (prep?)
- Short summary/synopsis of system test [pdf]
(Matt/Kurtis)
- Interleaved commissioning schedule (Gary)
- Gary (9/16 - 9/26) [Lynn Wood (9/18 - 9/25)]
- Kurtis (9/21 - 10/7) [Note: @ BAM 9/25 -
9/30]
- Matt A (9/23 - 10/2)
- Marc ?
- Preparations for bPID meeting on Thursday
- Electronics update/schedule (Kurtis/Matt)
- Update on beam test plans for Fermilab (Gary)
- Photonics assembly package, mounting plans/issues
(Marc)
- Review of
mechanical discussions at Nagoya (Gary)
- Review of further information regarding FTBF tables, iTOP
exoskeleton (Gary)
- Detailed design issues discussion (Marc, all
interested: must end by 3pm)
Follow-up/continuation
electro-mechanical discussion at 4pm as needed.
9-AUG-2011
Mechanical
meeting follow-up documentation
- Quartz Bar Box overview diagram [PNG]
- Quartz Bar Box assembly (dated 12-JUL-2010) [PDF]
- Inner Detector Support/TOP Module I/F (dated
27-AUG-2010) [PDF]
- Concept for seam strip bracket/Framing between QBBs
(10-AUG-2011) [PDF]
1-AUG-2011
Mechanics/Beam
Test Tracking meeting (2:00 - 3:00 pm)
- Short Review of the existing Kohriki CAD (Marc) [PDF]
- Revised what sent to Nagoya: [PDF]
- SciFi/hodoscope plans (Gary) [PDF]
- Beam test mechanical fixture (Marc) [PDF]
- AOB?
26-JUL-2011
Software
meeting (2:00 - 3:00 pm)
Dial-in information: (866) 740-1260 (toll free USA,
Canada)
Access code: 8147808 (followed by
"#")
- Further look at TDC distributions (Mike) [PDF] [Text]
- Documentation for Geant4 simulation (Matt B) [Text] (updated 2011/07/27)
[G4iTOP_20110727.tar.gz]
- Specific action items for next meeting?
19-JUL-2011
Software
follow-up meeting (2:00 - 3:00 pm)
- Further look at TDC distributions (Mike) [PDF]
- Comparison of MC of propagation time (Kurtis) [PDF]
- Additional slides (Matt B) [PDF]
- Questions from Gocha?
- Specific action items for next meeting?
18-JUL-2011
Local
mechanical/electrical
meeting (2:00 - 2:30 pm)
- Update on status of Nagoya system test preparations
(Kurtis/Matt A)
- Mechanics preparation for beam test (Marc)
- Discussion on ground rules for Wednesday's "FINAL REVIEW"
mechanics meeting with CDC (Marc, all) [PDF]
- SciFi development/discussion (Gary/Marc) --
separate discussion
Software/simulation/analysis
meeting
(2:30 - 3:00 pm)
Dial-in information: (866) 740-1260 (toll free USA,
Canada)
Access code: 8147808 (followed by
"#")
- Charges to this [PDF] group:
- Understand previous beam test data (match MC/data)
- Study reconstruction algorithms
- Optimize final detector configuration:
- mirror parameters (Yang)
- expansion wedge length, taper
- wavelength filter
- DOE review anticipated questions:
- Characterize Belle II performance with only dE/dx
PID?
- At what background level (random) do we get into
trouble?
- Sensitivity to "coherent" (temporal/spatial)
background photons?
- At what minimum photon level do we fall off a cliff
in terms of performance?
- Generate 2011 detector MC to guide beam test plan
- Cosmic ray MC/data check (BPAC - Nov. 11-12 timescale)
- Match detector MC/data from Fall 2011 beam tests (CD-1
review timescale -- early Feb. 2012)
- Update on Fall 2010 analysis
(Mike)
- Description [txt]
- And plots [PDF]
- Discussion on division of labor from dedicated PID
meeting in Nagoya (all) : detector-wide list [PDF]
- Update on GEANT4 simulation effort (Matt B) [PDF]
- Choice of time for regular version of this meeting? --> try
2pm (HST) on Tuesdays [5pm PDT, 8pm EDT]
- AOB?
July 4th, 2011
B2GM
postings
- Draft iTOP trigger summary slides [ppt]
(Gary for all)
- Update on beam test data analysis [pdf] (Mike)
20-JUN-2011
Local
meeting
(2:00 - 3:00 pm)
- CERN beam test data analysis status (Mike)
- Brief report on Simulation status and plans (Matt B)
- Update on board stack testing (Matt A)
- Electronics documentation update (Kurtis/Gary/all)
- System description document [link]
- BLAB3A datA sheet [link]
- SCROD firmware svn on google code [link]
- What can have
by B2GM/dedicated PID meeting (July 8-11)
- Full electronics module? (I leave Friday)
- Calibration optical fibers -- fiber part number?
- HV cables -- cable follow-up
- DC power cables -- choose cable
- Cooling pipes -- choose ID/OD
- Beam test additional info posted (Gary)
- Gantt Chart review (Kurtis)
- July and autumn scheduling plans discussion
- AOB?
31-MAY-2011
Local
meeting
(2:00 - 3:00 pm)
- CERN beam test data analysis status (Mike)
- Today's plots [PDF]
- Still no response from Arita-san to queries of 27-MAY
- Ongoing board designs:
- More SCROD boards?
- 3x ASIC motherboards
- Testing schedule before fab?
- Front boards
- HV board
- Design update Robin & Matt
- What can have
by B2GM/dedicated PID meeting?
- Full electronics module
- Calibration optical fibers -- fiber part number?
- HV cables -- choose cable
- DC power cables -- choose cable
- Cooling pipes -- choose ID/OD
- Gantt Chart Update
- July and autumn scheduling
- AOB?
23-MAY-2011
Local
meeting
(2:00 - 2:30 pm)
- Report from Japan (Tom)
- SL10's (16 channel) from Nagoya: 3x mock-up, 1x
functional
- Information on delivered PMTs (Inami-san [txt])
[PDF]
- Ongoing board design:
- 3x ASIC motherboards
- Testing schedule before fab?
- Front board
- Matt's study on geometry [pdf]
- HV board
- Carryover from previous discussion on mechanical mock-up:
- What can have by
B2GM/dedicated PID meeting?
- Calibration optical fibers -- fiber part number?
- HV cables -- choose cable
- DC power cables -- choose cable
- Cooling pipes -- choose ID/OD
- July and autumn scheduling
- CERN beam test data analysis status (Mike) [pdf]
- Update on discussion w/ Arita-san re: time resolutions
- AOB?
16-MAY-2011
Local
meeting
(2:00 - 2:30 pm)
- Report from Japan (Tom)
- information on delivered PMTs (Inami-san [txt]) [PDF]
- Update on mechanical mock-up -- what input still needed?
- What can have by
B2GM/dedicated PID meeting?
- Complete 3x other boards (Louie):
- Rest of carrier Stack (Levels 2-4) [layout done --
awaiting test]
- HV card [start soon]
- Front card [Matt to coordinate?]
- Calibration optical fibers -- fiber part number?
- HV cables -- choose cable
- DC power cables -- choose cable
- Cooling pipes -- choose ID/OD
- July and autumn scheduling [link]
- CERN beam test data analysis status (Mike)
- Reply from Arita-san [txt]
- Relevant plots [PDF]
- AOB?
9-MAY-2011
Local
meeting
(2:00 - 3:00 pm)
- Report from Japan (Tom)
- b-PID meeting this week
- update on Nov 2010 beam test analysis (Suzuki, Jones)
- progress towards autumn 2011 beam test:
- PMTs (Inami)
- electronics (Hawaii)
- mechanics (Hawaii)
- quartz (Horii)
- response from Kohriki-san to electronics mechanical
proposal (Suzuki)
- radiation test summary (Koga)
- quartz specifications confirmation (Inami, Shwartz)
- Plans for mechanical mock-up -- what still needed?
- Complete 3x other boards (Louie):
- Rest of carrier Stack (Levels 2-4) [layout done --
awaiting test]
- HV card [start soon]
- Front card [Matt to coordinate?]
- Calibration optical fibers -- choose fiber?
- HV cables -- choose cable
- DC power cables -- choose cable
- Cooling pipes -- choose ID/OD
- July and autumn scheduling
- CERN beam test data analysis status (from 2:30pm, when
Mike available)
- 9-MAY-2011 analysis update [link]
- AOB?
2-MAY-2011
Local
meeting
(2:00 - 3:00 pm)
- Updates on funding (Tom)
- Plans for mechanical mock-up -- what still needed?
- Complete 3x other boards (Louie):
- Rest of carrier Stack (Levels 2-4) [layout done --
awaiting test]
- HV card [start soon]
- Front card
- Calibration optical fibers -- choose component?
- HV cables -- choose cable
- DC power cables -- choose cable
- Cooling pipes -- choose ID/OD
- b-PID meeting this week? (golden week?)
- SCROD testing update
- Status of rest of cards
- Gantt update
- Anything new mechanics
- July and autumn scheduling
- CERN beam test data analysis status (from 2:30pm, when
Mike available)
- AOB?
18-APR-2011
Local
meeting
(2:00 - 3:00 pm)
- Review SuperBEAST WBS (Sven lead)
- spreadsheet [xlsx]
- Plans for mechanical mock-up -- what still needed?
(Marc lead discussion)
- Complete 3x other boards:
- Rest of carrier Stack (Levels 2-4)
- HV card
- Front card
- Calibration optical fibers?
- HV cables
- DC power cables
- Cooling pipes
- What to present at b-PID meeting this week?
- SCROD testing update
- Status of rest of cards
- Gantt update
- Anything new mechanics
- CERN beam test data analysis status (from 2:30pm, when
Mike available)
- Brief summary [txt]
- Further queries to Suzuki-san?
6-APR-2011
Board
Stack Design Review (10:30 am HST)
Cross-talk measured by Mike: details [txt] plot: [PDF]
Submission Review for 3 boards (SCROD-based readout):
- BLAB3A ASIC card
- Schematic [PDF]
- Layout [PDF]
- ASIC "carrier" card
- Schematic [PDF]
- Layout [PDF]
- Interface card
- Schematic [PDF]
- Layout [PDF]
- Submission schedule?
4-APR-2011
Local
meeting
(2:00 - 3:00 pm)
- Plans for mechanical mock-up? (Marc lead
discussion)
- Calibration optical fibers?
- HV cables
- DC power cables
- Cooling pipes
- Review WBS structure (1.3) and updates [4-APR-2011]
- Summary WBS spreadsheet (Jim Fast): Option
3: [xls]
Option 4: [xls]
- Readout Electronics (David Asner): Option
3:
[xls]
Option 4: [xls]
- B2GM meeting this week??
- CERN beam test data analysis status (from 2:30pm, when
Mike available)
- Fit results [PDF]
- Further queries to Suzuki-san (for which no response
received) [txt]
- Next time -- BEAST2 WBS review
30-MAR-2011
Special
WBS meeting (1:00 - 3:00 pm)
WBS Section 1.3 (Readout Electronics):
- US Belle II WBS overview diagram [PDF]
- Hawaii Labor rates basis: [xls]
- Summary WBS spreadsheet (Jim Fast): Option 3:
[xls] Option
4:
[xls]
- Readout Electronics (David Asner): Option
3:
[xls]
Option 4: [xls]
28-MAR-2011
Local
meeting
(2:00 - 3:00 pm)
Preparation for WBS Review on Wednesday is highest priority.
- Plans for mechanical mock-up? (Marc lead
discussion)
- Calibration optical fibers?
- HV cables
- DC power cables
- Cooling pipes
- Review WBS structure (1.3) and solicit electro-mechanical
input
- CERN beam test data analysis status (from 2:30pm, when
Mike available)
- Fit results [PDF]
- Further queries to Suzuki-san (for which no response
received) [txt]
- Continue and fill Section F (cost estimates -- continuing
later afternoon as needed), summary Spreadsheet
25-MAR-2011
Readout WBS meeting (1:00 - 3:00 pm)
WBS dedicated discussion in the IDL (Gary and Kurtis), with
the following reference documents:
- US Belle II WBS overview diagram [PDF]
- Summary WBS spreadsheet (Jim Fast) original: [xls]
IDL modified (option 3): [xls]
IDL modified (option 4): [xls]
- Readout Electronics (David Asner) original: [xls] IDL
modified (option 3): [xls]
IDL
modified
(option
4):
[xls]
BPAC Draft Report
(v2) [PDF]
21-MAR-2011
Local meeting (2:30 - 3:30 pm)
1) First pass at
Schedule/WBS
- Link to original file from Jim Fast: [xls]
- Discussion of items:
- Step 1:
- Getting task list correct
- Consider tasts that are specific elements of work and have
durations of 1 week to 2 months
- Level of effort tasks are generally not to be used
- Consider milestones as well. These will go in at 0
days duration and will be linked, generally, to completion
of some other task
- Step 2:
- Put appropriate resources on each task (at lowest level in
WBS).
- Equipment and materials should go in with dollars, e.g.
Equipment[5000].
- Labor should be listed with specific people names if
possible and then number of hours required to do the work on
the specific task
- Assignments:
- 1.2.1 PNNL
- 1.2.2 Alan (suspect will need a LOT of help)
- 1.2.3 Gary
- 1.2.4 Leo
- 1.2.5 Leo
- 1.2.6 Giovanni
- 1.2.7 Sven
- Modifications:
- Highlight new rows inserted and provide information
regarding predecessors and successors (i.e. how does this new
task fit in the flow of work)
- Upcoming meetings ? (mechanical this week?)
Confirm next B-PID?
2) Update on getting data sets from Nagoya? (Mike)
3) Follow-up from last PID meeting
on reflections at glue joints:
- Overall note on background photons in the DIRC: BaBar
Note 141 [PDF]
- Conclusions: scintillation unimportant, delta-ray
production dominates, with glue joint reflections of
second-most importance
- Underlying basis is measurements made in Note 140
- Specific measurements of reflection from Epotek joints:
BaBar Note 140 [PDF]
- Note wavelength dependence
- Non-Fresnel deviations seen
7-MAR-2011
Local meeting (2:30 - 3:30 pm)
1) Follow-up discussion from last mechanical meeting
- Marc on what is actual incursion? [PDF]
2) Update on getting data sets from Nagoya? (Mike)
3) Follow-up discussion from last PID meeting
- Fuji hall "clean room" clean enough for bar & bar-box
assy?
- Probably need space for electronics outside? (CR test
stand to be outside?)
- Stages and jigs?
- Gluing procedures? [coordination with Cincinnati?]
4) Revised (Inami-san) schedule for 2011 [PDF]
- Action item to revisit readout module preparedness Gantt Chart
- Commissioning with radiator should start end of July
- Several Belle II production (not R&D) SL-10 ready by
April
- 8 MCP-PMT ready by middle of April
- Concrete proposal for a CERN beam test [PDF]
- Started a [link] of items related to our preparations for this
[2011 iTOP beamtest]
5) Exhibits for the Ground Breaking Ceremony, next B2GM ??
(Discussion)
- needs to be before 3/18 ?? (for GBC)
- on exhibit in Tsukuba Exhibition after GBC?
- can hand-carry to April B2GM
6) DOE pow-wow March 15 in DC (no meeting next week?) -- additional
information needed? (Tom/Gary/Sven)
Follow-up
from
meeting: Kohriki drawing [PDF]
28-FEB-2011
Local meeting (Mike teaches until 2:30pm)
1) Reports from Belle II meeting
at
SLAC
- Marc on mechanics (post-SLAC meeting) [PDF]
- Critical issues -- DOE proposal/meeting March 15
- Schedule discussion (all)
- Revised cost estimate [xls]
2) Update on getting data sets from Nagoya? (Mike)
==> no response from Suzuki-san...
24-JAN-2011
No Regular meeting -- (Marc, Matt, Kurtis & Gary @ KEK)
1) A collection of pictures from the TOF dismount/repurpose (Jan
2011) [link]
2) Report from Yamaoka-san on the sag of the ECL (as measured on
flanges) [PDF]
3) Design Review at KEK (Jan 22, 2011) [txt summary] [IDM figure] [size XLS] [size PDF]
Conclusion is a request for 3mm
variance of the 10mm buffer zone.
Action
items:
- Screw intrusions into buffer zone to be requested from
Kohriki-san
- A more comprehensive cabling drawing is needed
4) Status at KEK:
- all TOF counters successfully moved to Fuji clean room area
- HV and cables being assembled
- First generation readout of 64 TOF channels possible with
CREAMTEA module (Matt will work on during his stay)
- Cut fiber optics need to be re-terminated; lengths to be
confirmed
- Geometry and support for the TOF counter configuration to be
optimized
5) Proposal by Inami-san for arrangement of the Fuji Hall
space [PDF]
17-JAN-2011
No Regular meeting -- MLK holiday (though many in for imminent
departure)
1) Report from Uehara-san at the Monday morning (KEK) meeting on TOF
extraction [PDF]
2) Report from Mike on studies of cosmic rays with Belle TOF,
indicating that 4cm RMS might be a realistic input value.
Data set: cosmic skim for Exp. 73, runs 849-916.
Posted file: [PDF]
Description: Plot 1
shows the counter number distribution and plot 2 shows the delta t distributions.
Because of the way the event T0 is determined, hits in one TOF counter will result in dT
near 0 and the dT in the
opposite counter will be near either -8 ns (top half) or 8 ns (bottom half of
detector).
Plots 3-4 show delta Z
versus zhit and the delta Z distribution respectively. For these plots delta Z is the calculated
z position (using the TDC
values for backward and forward PMTs) minus the zhit value
determined by extrapolation of the
track to the TOF.
One can see in plot 4
that the RMS is about 4 cm.
10-JAN-2011
Regular meeting
1) Report from
Marc on initial thoughts on TOF counter
configuration/support [PDF]
2) Initial
simulation studies of optimizing geometry/layout of Fuji Hall
CRT [PDF]
3) Schedules for
being at KEK:
Marc's schedule: Arrive KEK on the evening of Tue Jan 18th,
depart afternoon of Fri Jan 28th.
Gary:
arrive KEK evening of 21-JAN, depart NRT evening of 24-JAN
(25-26 JAN Trg/DAQ @ IHEP/Beijing, 28 JAN USTC/Hefei, 1-6
FEB Nanjing, 7-11 FEB SNU/Seoul)
Matt (KEK
support): Jan 20 - Feb. 4
Kurtis:
Jan 19 - Jan 24
4) Schedules
for rest of year -- Discussion
e.g. Gary to visit PNNL Friday; BGM/B2GM April
6-10
5) AOB?
27-DEC-2010
In lieu of meeting --
updated 31-DEC-2010:
Revised drawing (Marc) [jpg] and
zoomed-in dimensioned view [jpg]
--> First thoughts on bar configuration (Marc) [PDF]
==> Gary comments:
- can consider not close-packing the bars
- use 96 = 6 layers above and below, 8 bars/layer
- final geometry needs study
20-DEC-2010
Somewhat irRegular meeting (not all hands...)
1) Schedules for
being at KEK
Marc's schedule: Arrive KEK on the evening of Tue Jan 18th,
depart afternoon of Fri Jan 28th.
Gary:
arrive KEK evening of 21-JAN, depart NRT evening of 24-JAN
(25-26 JAN Trg/DAQ @ IHEP/Beijing, 28 JAN USTC/Hefei, 1-6
FEB Nanjing, 7-11 FEB SNU/Seoul)
Matt (KEK
support) -- dates
2) Cosmic test
stand design concept/plans
3) Response to Uehara-san's
message
4) discussion on
preparations and equipment for gluing -- visit to Zygo?
5) AOB?
13-DEC-2010
Regular meeting
1) Schedules for
being at KEK
Marc's schedule: Arrive KEK on the evening of Tue Jan 18th,
depart afternoon of Fri Jan 28th.
Gary:
arrive KEK evening of 21-JAN, depart NRT evening of 24-JAN
(25-26 JAN Trg/DAQ @ IHEP/Beijing, 28 JAN USTC/Hefei, 1-6
FEB Nanjing, 7-11 FEB SNU/Seoul)
Matt (KEK
support) -- dates? (Kurtis?)
2) Draft talk for PHENIX on DIRC detectors (Gary/Kurtis) (version 10) [PDF]
Draft talk for PHENIX on LAPPD/electronics
(Kurtis)
3) Schedules for rest of year
--> There will be a meeting
next week [link below]
4) AOB?
For next time: discussion on preparations and equipment for gluing --
visit to Zygo?
6-DEC-2010
Regular meeting
Follow-up
information (posted 7-DEC):
- on TOF Modules best for
experimenting upon (Mike Jones) [txt]
- Last set of HV bias
settings for the tubes (Uehara-san) [txt]
1) Updated schedule for TOF removal (from Uehara-san, for
discussion) [PDF]
2) Details on coupling of the suspension belts/brackets to TOF
mounts (Kichimi-san figure) [PDF]
3) Schedules for being at KEK?
Marc proposal: Arrive KEK on the evening of Tue Jan 18th,
depart afternoon of Fri Jan 28th.
Matt match?
(Kurtis?) Gary (en route to
IHEP/USTC/Korea....)
4) Plans for configuring the TOF bars/PMTs/readout. Including
fiber optic/laser calibration system. Reminder of existing bar
structure [PDF]
5) Straw-man proposal for a possible configuration (to be optimized
-- only 64 counters shown) [PDF]
6) How much of E-Hut electronics to keep??
Electronics/E-hut overview [PDF]
Itoh-san's message on clean-up rules
[txt]
Notes
from
Kichimi-san on instructions for dismounting/dismantling TOF
[PDF]
(posted
24-NOV-2010)
Follow-up video meeting on
Dec. 7th (Hawaii)/Dec. 8th KEK?
22-NOV-2010
Regular meeting
0) Belle II TDR available
1) Summary from Belle II B-PID [PDF] (Gary)
2) TOF dismounting plans discussion (all)
Message from Uehara-san:
Roll Out Dec.9
Removing cables Dec.28-30, (and also in January?)
(Extraction of CDC,BACC Jan.4-19)
Exraction of TOF Jan.20-26
Who, what, where, why and
when ??
3) Plans for cosmic test stand at Fuji ? (all)
4) Beam test at CERN? -- feedback (Gary)
5) 2011 preview of coming attractions (Gary)
Jan 25-26 B2 Trigger/DAQ in China
Feb ?? BAM/BAS
Apr 5-11 BGM/B2GM/Official commencement ceremony
6) AOB? (all)
8-NOV-2010
Regular meeting
1) Preparations for presentations next week [all]
a) readout update [Kurtis/Matt/Gary]
b) mechanics update (incl. electro-mech/board
stack) [Marc/Gary]
c) others?
2) Discussion of what to do once Zygo bars received? [all]
3) BLAB3 firmware update: status & plans [Kurtis/Matt/Gary]
4) Revised schedule for KEK/Nagoya installation/test trip (Feb/Mar 2011 ?) [Matt/Gary/Kurtis]
- dates for CERN beam test?
- existing or new board stack? (BLAB3 vs BLAB3A vs IRS2 w/amps)
- SL-10 PMT production test stand
5) AOB? [all]
25-OCT-2010
Regular meeting
0) 26-OCT-2010 -- Follow up on suggestion to check if other times 12x channels of ADC are zeros:
Thanks for the reply. I've checked ADC values for the other
modules and found
adc1 - 1206 events with all 12 ADC values =0
adc2 - 83 events with all 4 ADC values =0
adc3 - 13 events with all 12 ADC values =0
adc4 - 1 event with all 8 ADC values =0
It seems that whatever causes the zero ADC values is mostly on
the adc1 module.
If you make root files in the same format for the upcoming
CERN beam test, I would like to look at them.
1) Latest on Nagoya TOP counter analysis [Mike]
Message to Inami-san (21-OCT-2010):
Today I made a new nhit plot after modifying your script to
require for each channel
50<tdc[j] && tdc[j]<3000 && adc[j]>80 && adc[j]<500
where adc[j] is the ADC value for channel j.
The new plot is mdjnhitadclt500.ps. [PDF]
I added the adc[j]<500 requirement
because I had seen previously (plots c53.ps and c55.ps) that
some channels had a much broader distribution of adc values
for nhit=36 and nhit=24 without the adc[j]<500 requirement.
The mdjnhitadclt500.ps plot has nhit distributions with
veto (red histogram) and without veto (blue histogram) which
smoothly approach zero for nhit>30. In particular, there is
no peak at nhit=36. This could indicate that ADC values
above 500 are due to some kind of noise which somehow affects
all channels.
Further Message to Inami-san (22-OCT-2010):
Today I made a new nhit plot requiring for each channel
50<tdc[j] && tdc[j]<3000 && adc[j]>80 to calculate nhit.
The plot is mdjnhitadcanom.ps. [PDF]
The upper histogram is for events with at least one non-zero
ADC value among the 1st four channels
[(adc1_1+adc1_2+adc1_3+adc1_4)>0]. The lower histogram is for
1206 events for which (adc1_1+adc1_2+adc1_3+adc1_4)==0. The
lower histogram has no entries for nhit>24 because all 1206
events have the ADC values equal zero for the next eight
channels (adc1_5 through adc1_12). I checked one event
(evtnum==26) for which the 1st 12 channels have ADC values
equal zero and found that the TDC values look normal. Are
you aware of any anomaly which causes the 1st 12 channels
to have adc=0? Are these 12 channels read out with the same
electronics module?
Additional plots not sent (25-OCT-2010)
"event displays" for 4 events:
TDC vs. channel [run80ev3.ps] [PDF]
ADC vs. channel [run80ev3adc.ps] [PDF]
Event 26 is plotted with square symbols. The ADC values are zero
for channels 1-12.
2) Mechanics update [Marc]
a) beam test
b) mechanical mock-up
c) discussion on recent news [Tom/Gary]
3) Beam test at CERN/B2GM scheduling [Gary]
4) Overview of test cards/systems/new ID Lab system [Gary]
- [PDF]
5) BLAB3 firmware status & plans [Kurtis/Matt]
6) Revised schedule for KEK/Nagoya installation/test trip (beginning Jan 2011) [Matt/Gary/Kurtis]
7) AOB? [all]
18-OCT-2010
Regular meeting
1) Anything further on Nagoya TOP counter analysis [Mike]
Message to Inami-san (18-OCT-2010):
I made some plots to investigate the veto counter adc
distributions as a function of nhit. I used my modified script which requires
50<tdc[j] && tdc[j]<3000 && adc[j]>80
for each channel to calculate nhit. The plots c66.ps through
c70.ps show the veto adc versus nhit+evtnum/17000 with various cuts
on the number of veto counters with adc>80.
c66.ps veto counter 1 adc5_5>80||adc5_6>80 3853 entries [c66.pdf]
c67.ps veto counter 1 adc5_5>80 2631 entries [c67.pdf]
c68.ps veto counter 2 adc5_5>80||adc5_6>80 3853 entries [c68.pdf]
c69.ps veto counter 2 adc5_6>80 2399 entries [c69.pdf]
c71.ps veto counter 1 adc5_5>80&&adc5_6>80 1173 entries [c71.pdf]
c70.ps veto counter 2 adc5_5>80&&adc5_6>80 1173 entries [c70.pdf]
A common feature of all these plots is a clear excess at
adc>4000 for nhit=36 and a smaller excess at adc>4000 for nhit=24.
Do adc values > 4000 indicate very large pulse heights or do they indicate no data? If adc values > 4000 indicate
very large pulse heights, why is there an excess for nhit=24 but not for
nhit values between 25 and 35?
2) Mechanics update (?) [Marc] (will be away?)
3) Beam test at CERN? [Gary]
4) Overview of test cards/systems/new ID Lab system [Gary]
- [PDF]
5) BLAB3 firmware status & plans [Matt]
6) Revised schedule for KEK/Nagoya installation/test trip (beginning Jan 2011) [Matt/Gary/Kurtis]
7) AOB? [all]
11-OCT-2010
Regular meeting time
1) Further Update on cosmic ray data analysis of the Nagoya TOP counter prototype [Mike]
Response from Inami-san received. Made the change suggested
The resulting script is inami1005a.C in ~mdjbelle at KEK.
The plot I obtained using this script is [inamiringcolor.ps] [PDF]
I also modified this script so that I could apply a cut
if(50<tdc[j] && tdc[j]<3000 && adc[j]>80) h2->Fill(j,tdc[j]);
The modified script is inami1005b.C and the plot is [inamiringcuts.ps] [PDF]
I had noticed earlier that some channels have many fewer entries than typical channels and made some
plots of nhit comparing individual channels.
Plot [c61.ps] [PDF] compares the channel with tdc14_3 (upper histogram) with that
for channel with tdc14_1 (lower histogram).
Plot [c63.ps] [PDF] compares the channel with tdc13_3 (upper histogram) with that
for channel with tdc13_4 (lower histogram).
A large fraction of the 2106 entries for channel tdc13_4 are due to events with nhit=36.
2) Updates on the mechanics [Marc]
- beam components status
- mechanical mock-ups (required input/component ordering) [GSV/Louie will look into]
3) BLAB3 firmware status & plans [Matt]
4) Revised schedule for KEK/Nagoya installation/test trip (beginning Jan 2011) [Matt/Gary/Kurtis]
5) Feedback from Detector R&D Workshop [Gary]
6) Priority items? [all]
7) AOB? [all]
28-SEP-2010
Follow-up on TOP Cosmic Results:
To Inami-san: I made some plots to check the possible cross
talk which could produce the two narrow, parallel bands along the
diagonal of the TDC correlation plots which I made earlier.
For these studies I use two adjacent channels (tdc15_4 &
tdc15_5) in the PMT which shows the most prominent TDC correlation
and a different PMT (channels tdc14_3 & tdc14_4) with only a
faint TDC correlation.
- Plots c44.pdf and c45.pdf show the ADC values
versus tdc15_5-tdc15_4 for the two channels respectively.
- The analogous plots for the other PMT are c46.pdf and c47.pdf.
The striking difference is the concentration with ADC values below
80 in plots c44.pdf and c45.pdf which is not
present[ed] in the plots for the other PMT. This seems to
confirm that small ADC values are associated with most of the
anomalous TDC correlation.
Perhaps one should require ADC values greater than 80 for studies
using TDC values.
Subsequent plots to try and understand what is going on for these
events where all of the PMT channels are hit:
- Applying this cut: [mdjnhitadcgt80.pdf]
- For tdc15_5 (anomolous TDC correlation) [c48.pdf]
- For tdc14_14 (normal) [c49.pdf]
- Corresponding plots for ADC versus nhit+(evtnum/17000) for two
channels [c53.pdf] and
[c55.pdf]
27-SEP-2010
(back to regular meeting time of 3pm)
0) Any updates/news from Japan/KEK? [Tom/Sven]
1) Update on report on cosmic ray data analysis of the Nagoya TOP counter prototype [Mike]
2) Updates on the mechanics [Marc] slides [PDF]
- beam components status
- mechanical mock-ups (required input/component ordering) [GSV/Louie will look into]
3) BLAB3 firmware status & plans [Matt]
4) Revised schedule for KEK/Nagoya installation/test trip (beginning Jan 2011) [Matt/Gary/Kurtis]
5) Priority items? [all]
20-SEP-2010
(again, special short meeting -- 2:30-3pm)
With Tom, Sven at KEK and Kurtis remaining in an undisclosed location, the meeting should be relatively short,
do look forward to hearing from Mike on analysis of cosmic ray data, as well as status report from Marc about
the beam test stand mechanical preparations.
1) Report on cosmic ray data analysis of the Nagoya TOP counter prototype [Mike]
- Follow-up ADC plot for correlated events [PDF]
- Posted summary plots files [PDF]
2) Updates on the mechanics [Marc]
- beam components status
- mechanical mock-ups (required input/component ordering)
Marc had to go home due to sick kid, report via e-mail:
- BTF parts are all in shop queue. BTF commissioning, sans
QBB will be in Sven's lab W319.
Currently working on finalizing encoders. Plan is to ship
to Nagoya ASAP.
- No work on iTOP test stand,
but will bump to a higher priority.
- Plan to present new iTOP electro-mechanics slides @ pid_up
this week, to get feed back on
impact wrt CDC. Please
let
me know if there are any updates on our newest design
w/horizontal cards.
To be
discussed.
3) BLAB3 firmware plans [Matt]
4) Priority items? [all]
13-SEP-2010
(special short meeting -- 2:45pm)
1) Updates on the mechanics [Marc] [PDF]
- beam components status
- mechanical mock-ups (required input/component ordering)
2) BLAB3 firmware plans [Matt]
3) Priority items? [all]
7-SEP-2010
(archive)
1) No meeting this week due to intensive meetings last
2) Beam test preparations [all]
- post Kurtis' Gantt for weekly review
- available personnel
- status of expected performance
3) Updates on the mechanics [Marc]
- test beam configuration
- mechanical mock-ups
4) Plans for the next couple of months...
6) Visits:
- Matt 9/30 - 10/17 (& Mike with some overlap) thereafter
7) AOB?
Archival information:
List
of
dedicated
PID meeting (early September 2010) in Honolulu attendees
[link]
0) PID micro-group meeting from 3pm
1) Review of "final" detector configuration meeting (5/17) [Gary]
- what still remains?
- "2 weeks" to freeze
2) Final TDR revisions [based upon Review] (5/23 & 5/24)
Due 6/16!!
Uno/Haba-san comments [PDF] [docx]
3) Plans for full DOE proposal -- Due 7/1!! [Gary for Tom]
4) HI-TIDE HV/LV power options [Marc]
5) Summer schedule and discussion (building toward Nov beam test)
- mechanical mock-up
- BLAB3 module commissioning
- data processing development
6) AOB?
Deferred update items:
i) Final Simulation/reconstruction plans [Kurtis/all]
ii) mechanics/electronics packaging [Marc/???]
Reference Material:
For DOE proposal preparation (or perhaps later), a set of reference documents on costing
may be found in the /costing directory. A sample document for LBNE costing [PDF].
A note from Ushiroda-san on how to submit a Belle II note, with a direct [link]
There is a Trigger/DAQ meeting in between an earlier AMBER meeting and
the regular iTOP/bPID, which covers issues related to triggering and data
transfer/handling.
Mike Jones study of TS stretch factor/pedestal [Exp 57] [Exp 71]
Compiled TOF timing studies [Mike]
- Belle Note 1146 [PDF] [gs]
- Recent Timing studies [PDF] [gs]
- timing (T0) [Mike]
TDR {current} (Wiki repository)
Trigger/DAQ (after AMBER)
== Original TDR_v0.1 draft needs work TDR working page [all]
== Link to previous TDR source: [working directory]
== Message from Mike on timing studies with exiting TOF [20-FEB-2010/mdj]
Nov 2010 CERN beam test data [link]
further documentation on actual beamtest info [link]
Archive (snapshots):
Draft BPAC Beamtest/electronics talk [PDF]
Return to scheduling page
[UH Physics] [University of Hawaii]
Last modified: 2/14/2016
-- GSV